OpenStudy (anonymous):

Medal Simplify. (2/7)^2 - (- 2/7) ^2 = ?

OpenStudy (anonymous):

2^2/7^2- (-2)^2/7^2= 4/49-4/49= (4-4)/49=0

OpenStudy (anonymous):

You have to remember if the negative sign is inside the brackets (-2/7)^2 the whole expression becomes positive, so it will be 0 once you compute.

OpenStudy (anonymous):

so the answer is 0?

OpenStudy (anonymous):

yes 0

OpenStudy (anonymous):

thank you